OpenStudy (nsh4267):

Solve the inequality, and express the solutions in terms of intervals whenever possible. 4 different inequalities here.

OpenStudy (nsh4267):

a) \[x \left( 3x-1 \right)\le4\] b) \[\frac{ x ^{2}-x }{ x ^{2}+2x }\le0\] c) \[\frac{ x-2 }{ x^{2}-3x-10}\ge0\] d) \[\frac{ x+1 }{2x-3 }>2\]